# GraphIF: Enhancing Multi-Turn Instruction Following for Large Language Models with Relation Graph Prompt

### v1

@@ Introduction
Large Language Models (LLMs) (achiam2023gpt; dubey2024llama; team2024qwen2) have demonstrated exceptional performance in dialogue systems. As conversational AI becomes increasingly important, the ability to understand and follow user instructions is crucial for effective interaction (chiang2023vicuna; zheng2023judging). Real-world conversations pose a challenge as users typically distribute requirements across multiple dialogue turns, making multi-turn instruction following essential for maintaining dialogue coherence and consistency.
Recent efforts to enhance multi-turn instruction following have primarily focused on collecting (chiang2023vicuna; zhao2024wildchat) or generating (cao2025condor; gao2025strategic) multi-turn dialogue datasets to fine-tune LLMs. However, even well-tuned LLMs struggle with complex long-distance constraints (kwan2024mt). As illustrated in Figure 1, LLMs often forget instructions established in earlier turns, such as ending responses with specific phrases. This limitation stems from the mismatch between autoregressive training and the non-sequential dependencies in multi-turn dialogues—current instruction-tuning methods treat each response generation as an isolated task without explicitly modeling cross-turn constraints (zhang2025survey).
While incremental training could mitigate this issue, it incurs substantial computational costs and suffers from weak generalization due to data limitations (li2025beyond). This raises a fundamental question: can we develop training-free approaches that enable LLMs to satisfy inter-turn relational constraints without parameter updates?
@@
Graph structure has been proven to be able to represent the semantic relations in scenarios including multi-document QA (wang2024knowledge; edge2024local) and long-context QA (li2024graphreader). In multi-turn dialogues, the relational constraints between different dialogue turns can be naturally represented through labeled directed edges, making graph structures particularly suitable for modeling multi-turn instruction following (li2025structflowbench). Despite existing work on using graph structures to generate multi-turn dialogue datasets (li2025structflowbench), leveraging graph structures to enhance multi-turn instruction following remains unexplored.
@@
(1) We design an agent-based relation extraction module. Directly extracting semantic relations between dialogue turns is difficult, especially for complex long-distance constraints (bai2024mt; sun2024parrot). Inspired by agent-based decomposition paradigms (yao2023react; erdogan2025plan), we design an agent-based relation extraction module that employs LLM to alternate between action identification and action execution phases to iteratively construct the dialogue relation graph.
@@
Given that existing multi-turn instruction following datasets have limited dialogue turns and oversimplified inter-turn relations (zheng2023judging; he2024multi; bai2024mt), we construct two new datasets with extended dialogues and complex semantic relations based on two prior benchmarks (kwan2024mt; li2025structflowbench). Comprehensive experiments on the two datasets demonstrate that current fine-tuned LLMs exhibit suboptimal performance when confronted with complex constraints. Our GraphIF framework can be seamlessly integrated into existing LLMs and achieves significant improvements across all four multi-turn instruction-following evaluation metrics.
@@
Current approaches for enhancing multi-turn instruction following mainly use supervised fine-tuning with instruction datasets. Some methods curate real-world user-LLM interactions (wang2023openchat; zhao2024wildchat), while others generate synthetic dialogues (ding2023enhancing; wu2025instruct). Parrot (sun2024parrot) addresses anaphora and ellipsis by training specialized models, and ConsistentChat (chen2025consistentchat) tackles consistency through skeleton-guided generation. However, these instruction-tuning methods overlook explicit modeling of inter-turn relational structures.
@@
Recent research has increasingly explored the integration of graph structures to enhance the capabilities of LLMs (jimenez2024hipporag; li2024graphreader; lin2025rje). GraphRAG (edge2024local) and LightRAG (guo2024lightrag) construct cross-document knowledge graphs to enable knowledge-enhanced retrieval for response generation. KGP (wang2024knowledge) addresses multi-document question-answering tasks by constructing passage-level directed graphs to aggregate relevant information. However, no existing research has applied graph structures to enhance the multi-turn instruction following for LLMs.
@@
Given complex relational dependencies between current instructions and multiple dialogue turns, single-step relation extraction is insufficient. Inspired by agent-based problem decomposition paradigms (yao2023react; erdogan2025plan), we propose an iterative agent-based module for progressive relation extraction. Since different dialogue relations require tailored identification approaches, our module decomposes relation extraction into two complementary subtasks: Action Identification and Action Execution.
@@
We define a set of actions that characterize representative dialogue relations. Unlike previous benchmarks (kwan2024mt; li2025structflowbench), we define a more inclusive and coarse-grained set of relations that better capture the dialogue constraints in real-world scenarios.
@@
Qwen2.5-3B
-Instruct
LLM-Only 59.42 9.56 59.09 56.99
@@
Qwen2.5-7B
-Instruct
LLM-Only 80.22 51.30 79.70 79.57
@@
Qwen2.5-14B
-Instruct
LLM-Only 90.22 75.65 90.00 92.61
@@
Existing multi-turn instruction-following datasets are limited by short dialogue turns (typically under eight turns) (bai2024mt; he2024multi), inadequately representing complex multi-turn instruction following scenarios. To address this, we construct evaluation datasets following two benchmarks that model inter-turn relations. We design MT-Eval* by merging and manually verifying multiple instances from MT-Eval (kwan2024mt). Similarly, we leverage the customizable structural framework of StructFlowBench (li2025structflowbench) to design StructFlowBench*. The two datasets we construct feature dialogues with over 20 turns, each containing multiple inter-turn relations within individual dialogue instances.
@@
We evaluate five instruction-tuned LLMs as backbone architectures, comprising two official models: Llama-3.1-8B-Instruct (dubey2024llama) and Qwen2.5-7B-Instruct (team2024qwen2), along with three models that are fine-tuned and aligned through preference-based methods based on Llama-3.1-8B: Hermes-3-Llama-3.1-8B [^1] https://huggingface.co/NousResearch/Hermes-3-Llama-3.1-8B, Llama-3.1-Storm-8B [^2] https://huggingface.co/akjindal53244/Llama-3.1-Storm-8B, and Llama-3.1-Tulu-3.1-8B [^3] https://huggingface.co/allenai/Llama-3.1-Tulu-3.1-8B (lambert2024tulu).
@@
  • MemoryBank (zhong2024memorybank) summarizes dialogue content as events and employs RAG mechanisms to retrieve relevant conversations, enhancing models’ memory of key information.
  • MemoChat (lu2023memochat) creates topic-indexed storage structures and summarizes related dialogues, using LLM-based retrieval to enhance response generation.
@@
We adopt four instruction-following metrics to evaluate models from different perspectives: Constraint Satisfaction Rate (CSR), Instruction Satisfaction Rate (ISR) (zhang2024cfbench), Decomposed Requirements Following Ratio (DRFR) (qin2024infobench), and Weighted Constraint Satisfaction Rate (WCSR) (li2025structflowbench). Following recent benchmarks (zheng2023judging; li2025structflowbench), we utilize GPT-4o as the LLM judge with human verification.
@@ Acknowledgments
Qwen2.5-3B
-Instruct
LLM-Only 67.42 20.23 67.88 63.61
@@
Qwen2.5-7B
-Instruct
LLM-Only 70.70 23.02 71.22 67.07
@@
Qwen2.5-14B
-Instruct
LLM-Only 74.49 31.64 75.36 71.09
@@
Qwen2.5-7B
-Instruct
LLM-Only 1 3.01 1 7.78
@@
Qwen2.5-14B
-Instruct
LLM-Only 17.70 10.00 18.84 21.95 17.12

### v2

@@ Introduction
Large Language Models (LLMs) (Achiam et al. 2023; Dubey et al. 2024; Team 2024) have demonstrated exceptional performance in dialogue systems. As conversational AI becomes increasingly important, the ability to understand and follow user instructions is crucial for effective interaction (Chiang et al. 2023; Zheng et al. 2023). Real-world conversations pose a challenge as users typically distribute requirements across multiple dialogue turns, making multi-turn instruction following essential for maintaining dialogue coherence and consistency.
Recent efforts to enhance multi-turn instruction following have primarily focused on collecting (Chiang et al. 2023; Zhao et al. 2024) or generating (Cao et al. 2025; Gao et al. 2025) multi-turn dialogue datasets to fine-tune LLMs. However, even well-tuned LLMs struggle with complex long-distance constraints (Kwan et al. 2024). As illustrated in Figure 1, LLMs often forget instructions established in earlier turns, such as ending responses with specific phrases. This limitation stems from the mismatch between autoregressive training and the non-sequential dependencies in multi-turn dialogues—current instruction-tuning methods treat each response generation as an isolated task without explicitly modeling cross-turn constraints (Zhang et al. 2025).
While incremental training could mitigate this issue, it incurs substantial computational costs and suffers from weak generalization due to data limitations (Li et al. 2025b). This raises a fundamental question: can we develop training-free approaches that enable LLMs to satisfy inter-turn relational constraints without parameter updates?
@@
Graph structure has been proven to be able to represent the semantic relations in scenarios including multi-document QA (Wang et al. 2024; Edge et al. 2024) and long-context QA (Li et al. 2024). In multi-turn dialogues, the relational constraints between different dialogue turns can be naturally represented through labeled directed edges, making graph structures particularly suitable for modeling multi-turn instruction following (Li et al. 2025a). Despite existing work on using graph structures to generate multi-turn dialogue datasets (Li et al. 2025a), leveraging graph structures to enhance multi-turn instruction following remains unexplored.
@@
(1) We design an agent-based relation extraction module. Directly extracting semantic relations between dialogue turns is difficult, especially for complex long-distance constraints (Bai et al. 2024; Sun et al. 2024). Inspired by agent-based decomposition paradigms (Yao et al. 2023; Erdogan et al. 2025), we design an agent-based relation extraction module that employs LLM to alternate between action identification and action execution phases to iteratively construct the dialogue relation graph.
@@
Given that existing multi-turn instruction following datasets have limited dialogue turns and oversimplified inter-turn relations (Zheng et al. 2023; He et al. 2024; Bai et al. 2024), we construct two new datasets with extended dialogues and complex semantic relations based on two prior benchmarks (Kwan et al. 2024; Li et al. 2025a). Comprehensive experiments on the two datasets demonstrate that current fine-tuned LLMs exhibit suboptimal performance when confronted with complex constraints. Our GraphIF framework can be seamlessly integrated into existing LLMs and achieves significant improvements across all four multi-turn instruction-following evaluation metrics.
@@
Current approaches for enhancing multi-turn instruction following mainly use supervised fine-tuning with instruction datasets. Some methods curate real-world user-LLM interactions (Wang et al. 2023; Zhao et al. 2024), while others generate synthetic dialogues (Ding et al. 2023; Wu et al. 2025). Parrot (Sun et al. 2024) addresses anaphora and ellipsis by training specialized models, and ConsistentChat (Chen et al. 2025) tackles consistency through skeleton-guided generation. However, these instruction-tuning methods overlook explicit modeling of inter-turn relational structures.
@@
Recent research has increasingly explored the integration of graph structures to enhance the capabilities of LLMs (Jimenez Gutierrez et al. 2024; Li et al. 2024; Lin et al. 2025). GraphRAG (Edge et al. 2024) and LightRAG (Guo et al. 2024) construct cross-document knowledge graphs to enable knowledge-enhanced retrieval for response generation. KGP (Wang et al. 2024) addresses multi-document question-answering tasks by constructing passage-level directed graphs to aggregate relevant information. However, no existing research has applied graph structures to enhance the multi-turn instruction following for LLMs.
@@
Given complex relational dependencies between current instructions and multiple dialogue turns, single-step relation extraction is insufficient. Inspired by agent-based problem decomposition paradigms (Yao et al. 2023; Erdogan et al. 2025), we propose an iterative agent-based module for progressive relation extraction. Since different dialogue relations require tailored identification approaches, our module decomposes relation extraction into two complementary subtasks: Action Identification and Action Execution.
@@
We define a set of actions that characterize representative dialogue relations. Unlike previous benchmarks (Kwan et al. 2024; Li et al. 2025a), we define a more inclusive and coarse-grained set of relations that better capture the dialogue constraints in real-world scenarios.
@@
Qwen2.5-3B -Instruct LLM-Only 59.42 9.56 59.09 56.99
@@
Qwen2.5-7B -Instruct LLM-Only 80.22 51.30 79.70 79.57
@@
Qwen2.5-14B -Instruct LLM-Only 90.22 75.65 90.00 92.61
@@
Existing multi-turn instruction-following datasets are limited by short dialogue turns (typically under eight turns) (Bai et al. 2024; He et al. 2024), inadequately representing complex multi-turn instruction following scenarios. To address this, we construct evaluation datasets following two benchmarks that model inter-turn relations. We design MT-Eval* by merging and manually verifying multiple instances from MT-Eval (Kwan et al. 2024). Similarly, we leverage the customizable structural framework of StructFlowBench (Li et al. 2025a) to design StructFlowBench*. The two datasets we construct feature dialogues with over 20 turns, each containing multiple inter-turn relations within individual dialogue instances.
@@
We evaluate five instruction-tuned LLMs as backbone architectures, comprising two official models: Llama-3.1-8B-Instruct (Dubey et al. 2024) and Qwen2.5-7B-Instruct (Team 2024), along with three models that are fine-tuned and aligned through preference-based methods based on Llama-3.1-8B: Hermes-3-Llama-3.1-8B [^1] https://huggingface.co/NousResearch/Hermes-3-Llama-3.1-8B, Llama-3.1-Storm-8B [^2] https://huggingface.co/akjindal53244/Llama-3.1-Storm-8B, and Llama-3.1-Tulu-3.1-8B [^3] https://huggingface.co/allenai/Llama-3.1-Tulu-3.1-8B (Lambert et al. 2024).
@@
  • MemoryBank (Zhong et al. 2024) summarizes dialogue content as events and employs RAG mechanisms to retrieve relevant conversations, enhancing models’ memory of key information.
  • MemoChat (Lu et al. 2023) creates topic-indexed storage structures and summarizes related dialogues, using LLM-based retrieval to enhance response generation.
@@
We adopt four instruction-following metrics to evaluate models from different perspectives: Constraint Satisfaction Rate (CSR), Instruction Satisfaction Rate (ISR) (Zhang et al. 2024), Decomposed Requirements Following Ratio (DRFR) (Qin et al. 2024), and Weighted Constraint Satisfaction Rate (WCSR) (Li et al. 2025a). Following recent benchmarks (Zheng et al. 2023; Li et al. 2025a), we utilize GPT-4o as the LLM judge with human verification.
@@ Acknowledgments
Qwen2.5-3B -Instruct LLM-Only 67.42 20.23 67.88 63.61
@@
Qwen2.5-7B -Instruct LLM-Only 70.70 23.02 71.22 67.07
@@
Qwen2.5-14B -Instruct LLM-Only 74.49 31.64 75.36 71.09
@@
Qwen2.5-7B -Instruct LLM-Only 1 3.01 1 7.78
+GraphIF 4.94 10.89 4.96 19.56
Llama-3.1-8B -Instruct LLM-Only 1 3.01 1 7.78
@@
Qwen2.5-14B -Instruct LLM-Only 17.70 10.00 18.84 21.95 17.12